OpenStudy (anonymous):

simpliy [(3+5)*2-5]*5

OpenStudy (bloomlocke367):

Welcome to OpenStudy!!! With this question, you just have to follow the order of operations. Are you familiar with that?

OpenStudy (anonymous):

yes

OpenStudy (bloomlocke367):

Knowing that, what part do you think you'd do first?

OpenStudy (anonymous):

but the only answers i can get are a9 b19 c49 d29 but I get 33

OpenStudy (tkhunny):

You should answer the question that was presented by Bloom.

OpenStudy (anonymous):

whats the answer

OpenStudy (bloomlocke367):

I can't just give you answers, it's against the Code of Conduct.

OpenStudy (bloomlocke367):

But, I CAN help you reach the answer

OpenStudy (anonymous):

ok

OpenStudy (bloomlocke367):

Now, which part do you think you do first, using the order of operations (PEMDAS)?

OpenStudy (bloomlocke367):

It's okay if you have no idea, you just have to let me know so I can help you.

OpenStudy (anonymous):

parentheses witch is 8. then 8 * 2 = 16, 16 -5 = 11, finally 11* 3 = 33. witch is not an answer.

OpenStudy (tkhunny):

Is it a 3 or a 5 or something else on the end?

OpenStudy (bloomlocke367):

I was just about to ask that @tkhunny XD

OpenStudy (tkhunny):

Sorry about that. I get a little jumpy, on occasion.

OpenStudy (bloomlocke367):

That's fine :D I don't mind you helping too, lol. I just thought it was funny we had the same thought.

OpenStudy (bloomlocke367):

Welp, I have to go now. Best of luck @mathmaster666! I hope you figure it out soon :D

OpenStudy (tkhunny):

Sometimes, answer are wrong. Sometimes, problems are printed incorrectly. Go with confidence. You seem to have a clue.

OpenStudy (bloomlocke367):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @tkhunny Sometimes, answer are wrong. Sometimes, problems are printed incorrectly. Go with confidence. You seem to have a clue. \(\color{#0cbb34}{\text{End of Quote}}\) Nicely said! :D